Please describe me about wheatstone bridge, its applications and conditions where it can and can't be used.


I will very very thankful to you

Grade:

2 Answers

DIVYA SHARMA
11 years ago

OK WHEATSTONE BRIDGE IS NOTHING BUT JUST A CIRCUIT WHICH FOLLOWS KIRCHOFFS LOOP RULE,

I WILL EXPLAIN U ON BASIC OF NCERT DIAGRAM.

JUST BEFORE I EXPLAIN U TAKE THIS BASIC CONVENTION ALSO

JUNCTION RULE:SUM OF CURRENT ENTERING A JUNCTION=SUM OF CURRENT LEAVING.

NOW SEE NCERT DIAGRAM:

I1=I3,I2=I4(NO CURRENT GOES THROUGHT GALV. ARM)

IN CLOSED LOOP ABDA:

V2-V1=0(LOOP RULE)(AT JUNCTION AS CURRENT IS NEGATIVE POTENTIAL IS NEGATIVE)

HEANCE I2R2 - I1R1=0 =>I2R2=I1R1........(1)

SIMILARLY IN CLOSED LOOP BCDB:

I2R4=I1R3......(2)

DIVIDE THEM U GET (R1/R2 = R3/R4)

THIS IS WHEATSTONE BRIDGE(MAIL AGAIN IF U NEEDED ITS AIM AND CONS.) APPLICATION:METRE BRIDGE APPLICATIONS

DIVYA SHARMA
11 years ago

SORRY FOR LAST LINES:

APPLICATION OF WHEATSTONE BRIDGE LIES IN METER BRIDGE AN POTENTIOMETER

U WILL DO THESE IN UR PRACTICALS ALSO:

METER BRIDGE:

1. TO FIND RESISTIVITY OF GIVEN MATERIAL.

2. TO FIND VALUE OF UNKNOWN RESISTANSE USING METER BRIDGE.

POTENTIOMETER:

1.TO FIND OUT EMF OF AN UNKNOWN BATTERY USING 1 KNOWN BATTERY

HOPE U ARE CLEAR NOW